# Distributed Maximum Flow in Planar Graphs

### v3

@@ 6.2 Minimum st-Cut
Given a directed planar graph $G$ with non-negative edge weights and two vertices $s,t$, there is a $\tilde{O}(D^{2})$ round randomized algorithm that finds the minimum $st$-cut w.h.p.
To prove this, we use Theorem 1.2 to compute a maximum $st$-flow $f$ in $G$ in $\tilde{O}(D^{2})$ rounds. By [7], a corresponding minimum $st$-cut is the cut defined by the saturated edges (an edge is saturated if the flow it pushes equals its capacity) that precede any other saturated edge on any path that contains them and starts at $s$. To find those edges we set the weight of each saturated edge to one and all other edges to zero. Then, we use an exact distance algorithm that runs in $\tilde{O}(D^{2})$ rounds for planar graphs by [26] on the graph with modified weights. Each edge that has an endpoint with distance zero from $s$ and another endpoint of distance one from $s$ is a cut edge. The bisection can be deduced from the distances as well, each vertex whose distance from $s$ is zero on $s$’s side of the cut, others are on $t$’s side. For a full proof, the reader is referred to Section D.2.
Next, we show an almost optimal algorithm for $(1+\epsilon)$ approximation of the undirected minimum $st$-cut, where $s$ and $t$ lie on the same face.
@@
Given an undirected planar graph $G$ with non-negative edge weights and two vertices $s,t$ that lie on the same face of $G$, there is a randomized algorithm that finds an approximate $(1+\epsilon)$ minimum $st$-cut and runs in $D\cdot n^{o(1)}$ rounds w.h.p.
@@
Concretely, we take a primal-dual detour, where we consider faces in the dual graph that map to the vertices $s$ and $t$ in the primal [^12] The dual of the dual is the primal, i.e, primal vertices map to faces of the dual graph. Moreover, the dual face that is mapped to any vertex $v$ is the (only) face that consists of all darts entering $v$. See [25] for more details.. See Figure 14. We refer to faces of the dual as dual faces. A special case of the cycle-cut duality (3.1) is given by Reif [38], stating: An $st$-separating cycle in the dual graph is the dual of an $st$-cut in the primal, where a cycle in the dual graph is said to be $st$-separating, iff it contains the the dual face that is mapped to $s$ in its interior, and the dual face mapped to $t$ in its exterior (or vice versa). Thus, we find a $(1+\epsilon)$-approximation of the minimum $st$-separating cycle in $G^{*}$ in order to find its dual, the desired $(1+\epsilon)$-approximate minimum $st$-cut in $G$.
@@
Note, any $st$-cut in $G\cup\{e\}$ shall contain $e$, and removing it gives $G$. Thus, to have the min $st$-cut in $G$, we can look for the min $st$-cut in $G\cup\{e\}$. In particular, there is a mapping between $st$-cuts of $G$ and $G\cup\{e\}$, where the value of the cut in $G\cup\{e\}$ is the same value of the corresponding cut in $G$ plus the weight ($nW$) of $e$. Thus, when applying the duality of Reif we consider $G^{*}_{virt}$. That is, we look for the approximate minimum separating cycle in $G^{*}_{virt}$. By duality, this cycle should contain $e^{*}=(f_{1},f_{2})$.
@@
Note, the edge $e^{*}=(f_{1},f_{2})$ closes a cycle with $P$, denoted $C=P\cup\{e^{*}\}$, that would be the desired approximate min $st$ separating cycle. $C$ is separating as any cycle in a planar graph is separating (Jordan curve theorem [44]), in particular it is $st$ separating, for: (1) $s$ and $t$ are faces in the dual, each has to be either in the interior or the exterior of $C$, (2) one dart of $e$ is in $C$’s interior and the other in its exterior, each such dart corresponds to a distinct dual face, one of which maps to $s$ and the other maps to $t$ because $e$ connects $s$ and $t$ in the primal. Note, $C$ is $(1+\epsilon)$-approximate $st$-separating minimum cycle in $G^{*}_{virt}$. Otherwise, $P$ is not a $(1+\epsilon)$-approximation to the shortest $f_{1}$-to-$f_{2}$ path. I.e, since $C=P\cup\{e\}$, then the length of $C$ is that of $P$ plus $nW$. I.e, if $C$ is not a $(1+\epsilon)$ approximation of the minimum $st$ separating cycle, then, the length of $C$ minus $nW$ is not a $(1+\epsilon)$ approximation to the shortest $f_{1}$-to-$f_{2}$ path, contradicting the fact that $P$ is.
@@
Next, we aim to mark those $st$-cut edges in $G$. By the above, it is sufficient to mark edges in $G$ who are the duals of edges $P$ in $G^{*}_{virt}$. Assuming we already applied Theorem 1.3, in particular, we have computed an approximate SSSP tree $T$ from $f_{1}$, then, we want to mark $P$. We already know that we can simulate any minor-aggregation algorithm on $G^{*}_{virt}$ (for details, one is referred to the proof of Theorem 1.3), thus, we can simulate any minor-aggregation algorithm on any subgraph of $G^{*}_{virt}$ (Corollary 11 of [18]). I.e., we can invoke any minor-aggregation algorithm on $T$. To mark $P$ edges, a minor-aggregation tree-rooting and subtree-sum procedures of [18] are invoked, where we root $T$ with $f_{1}$ and compute subtree-sum on the rooted tree, such that, the input value of all nodes is zero, except for $f_{2}$ that would have an input of one. The edges whose both endpoints have subtree-sum value of one are marked, those exactly are the $P$-edges. Those procedures are of $\tilde{O}(1)$ minor-aggregation rounds, which are simulated within $\tilde{O}(D)$ rounds in $G$ by applying Theorem B.14, the exact details of this simulation on $G^{*}_{virt}$ are provided in the proof of Theorem 1.3.
@@
Next, we show how to find the cut vertices, in particular, we show how to find a set $S$ such that $P=\delta(S,V\setminus S)$. To do that, we take an approach similar to that described in the proof of Theorem 6.1, where we define a graph $R^{\prime}$, in which, each $P$-edge is assigned a weight of one, and all other edges a weight of zero, we then again use the minor-aggregation SSSP approximation algorithm of [42] with an arbitrary constant approximation parameter $\epsilon$ in order to compute a $(1+\epsilon)$-approximate SSSP tree $T$ from $s$ in $R^{\prime}$. In particular, we modify the algorithm to work with edges with weights in $[0,n^{O(1)}]$ rather in $[1,n^{O(1)}]$, as shown in the proof of Theorem 1.3. Finally, vertices whose distance is zero from $s$ are in $S$ (other vertices are in $V\setminus S$).
@@
The round complexity of simulating the SSSP approximation algorithm is $D\cdot n^{o(1)}$, since the part-wise aggregation problem can be solved on $G$ in $\tilde{O}(D)$ $\mathsf{CONGEST}$ rounds (Corollary B.6), and the approximate SSSP algorithm of [42] runs in $n^{o(1)}$ minor-aggregation rounds. Thus, by Lemma B.8 simulating the algorithm incurs a total of $D\cdot n^{o(1)}$ $\mathsf{CONGEST}$ rounds in $G$. Note, after applying the distance approximation algorithm, vertices of $G$ know locally to what side of the cut do they belong. ∎
@@ 7 Directed Global Minimum Cut
Preprocessing: We work with the network obtained from $G$ by adding the reversal dart of each edge to the graph, such that, the two darts of an edge are ordered one after another in the embedding. We abuse notation and denote this network by $G$. The reversal dart gets assigned a weight of zero. It is not hard to see that the new network can be constructed and embedded in the plane without any need for communication (assuming an embedding of the original network). Note that any $\mathsf{CONGEST}$ round on the augmented network is simulated within two rounds on the communication graph.
@@
    $ID(C_{X_{i}})$ for each child bag $X_{i}$ of $X$ is aggregated on $X$ in $\tilde{O}(D)$ rounds so that the ID of the minimum weight cycle $C_{X}$ (that does not cross $F_{\!{}_{X}}$) is known to all vertices in $X$. Note, vertices in $X_{i}$s know $ID(C_{i})$, thus they can broadcast it (this is obviously true for leaf bags, for non-leaf bags, this is recursively maintained in the next two steps). Since $X$ is of $\tilde{O}(D)$ diameter, this aggregation terminates in $\tilde{O}(D)$ rounds. Next, we consider cycles that cross $F_{\!{}_{X}}$.
@@
      Due to the above claim, we get that $P$ combined with $d$ gives the minimum weight simple dart cycle that uses $d$, denoted $C_{d}$. We define $C$ to be the minimum over $C_{d}$ for all $S_{\!{}_{X}}$ darts $d$. If $C$ is of weight strictly less than $C_{X}$, then $C_{X}$ is set to $C$ and is identified with the ID of $f$ concatenated with the cycle’s weight.
@@
      $C$ does not use $S_{\!{}_{X}}$: In this case, we iterate over node-parts $g_{i}$ in $F_{\!{}_{X}}$, in each iteration we remove $S_{\!{}_{X}}$ darts from $DDG(f)$ and all clique edges between all node-parts $g_{2},g_{3},\ldots$ that correspond to the same face $g$ of $G$ as $g_{i}$.
      Figure 17: $DDG(f)$ in the iteration of node-part $g_{2}$. $g_{1},g_{2},g_{3}$ correspond to the face $g$ of $G$, where each shadow color demonstrates a distinct child bag of $X^{*}$. Black vertices are endpoints of $S_{\!{}_{X}}$ edges and red vertices are node parts in $X^{*}$’s child bags. All $S_{\!{}_{X}}$ darts are removed from the DDG, as well as clique edges on $g_{1},g_{2},g_{3}$. All other edges have both their darts present (for simplicity, not some are undirected in the picture). The green directed path demonstrates a shortest $g_{2}$-to-$g_{1}$ path, representing a cycle that contains $g$ in $G^{*}$.
@@
    Note that $T$ is represented by $\hat{G}$, our tool for aggregations on $G^{*}$. Thus, we next work with it and with minor-aggregations on $G^{*}$ and $T$, which by Section 4 is possible since (1) we can do that on $G^{*}$ or any minor of it, and (2) $T$ is a subgraph (minor) of $G^{*}$. For an extended discussion of $\hat{G}$ and minor-aggregations on $G^{*}$, see Sections 3 and 4, and Appendices A and B.
@@
Appendix A The Graph $\hat{G}$
We next give the full definition of the graph $\hat{G}$, and prove its properties. Let $\deg(v)$ denote the undirected degree of a vertex $v$ in $G$. Then, a vertex $v$ in $G$ has $\deg(v)+1$ copies in $\hat{G}$: A copy $v$ referred to as a star center (we denote by $V_{S}=V(G)$ the set of all star centers) and a copy $v_{i}$ for each local region of $v$. A local region of $v$ is defined by two consecutive edges in $v$’s local embedding. Notice that $v$ has $\deg(v)$ local regions, and that each local region belongs to a face of $G$ (it is possible that several local regions of $v$ belong to the same face). Eventually, we will have that the faces of $G$ correspond to (vertex and edge disjoint) faces of $\hat{G}$. If $i$ local regions of $v$ belong to the same face $f$ of $G$, then $v$ will have $i$ copies in the face of $\hat{G}$ that corresponds to $f$. The vertices of $\hat{G}$ are therefore $V_{S}\cup_{v\in G}N(v)$ where $N(v)=\{v_{1},\ldots,v_{\deg(v)}\}$.
Figure 18: On the left, the primal graph $G$. The clockwise local numbering of $u$ and $v$’s edges is in red. $f_{\infty}$ denotes the infinite face of $G$. On the right, the face-disjoint graph $\hat{G}$. The face $f_{\infty}$ of $G$ maps to the face $C_{\infty}$ (thick black) of $\hat{G}[E_{R}]$. Vertex $v$ in $G$ has a star-center copy $v$ in $V_{s}\subset V(\hat{G})$, and $\deg(v)$ additional duplicates $v_{0},v_{1},v_{2}$ in $V(\hat{G})$ connected by star edges $E_{S}\subset E(\hat{G})$ (dashed blue). The edge $(u,v)$ of $G$ maps to three edges in $\hat{G}$: an edge $(v_{0},v_{2})$ in $E_{C}\subset E(\hat{G})$ (solid blue) and two edges $(u_{2},v_{2})$ and $(u_{1},v_{0})$ in $E_{R}\subset E(\hat{G})$ (black edges). The vertex $w\in G$ is of degree one, thus, it has a star-center copy and one additional copy in $\hat{G}$ connected by a dashed blue edge. Its only incident edge in $G$ has two duplicates, each connected to a distinct non star-center copy of its only neighbor. The non star-center copies of $w$’s neighbor are connected by a solid blue edge (corresponding to the edge incident to $w$), this edge corresponds to a self loop in $G^{*}$ as it is connects the same face to itself (the face containing $w$).
$E_{S}$:
These edges connect every star center $v$ and its copies $v_{i}\in N(v)$.
$E_{R}$:
For each edge $(u,v)\in G$, there are two edges in $E_{R}$, one copy for each face of $G$ that contains $(u,v)$. Let $(u,v)$ be the $i$’th and $j$’th edge in the clockwise local embedding of $u$ and $v$, respectively. Then the edges added to $E_{R}$ are $(u_{i+1},v_{j})$ and $(u_{i},v_{j+1})$. This is because the face containing the $i$’th and $(i+1)$’th edges of $u$ is the face containing the $j$’th and $(j-1)$’th edges of $v$ (and the face containing the $i$’th and $(i-1)$’th edges of $u$ is the face containing the $j$’th and $(j+1)$’th edges of $v$).
$E_{C}$:
In order to simulate computations on $G^{*}$, we make sure that if two faces of $G$ share an edge $e$ (i.e. are adjacent nodes in $G^{*}$) then their corresponding faces in $\hat{G}$ are connected by a corresponding edge in $E_{C}$, between two vertices $v_{j},v_{j+1}$ of the same star, since $v_{j},v_{j+1}$ each corresponds to a face that contains $e$. For an edge $(u,v)$ we can add an edge in $E_{C}$ between copies of $u$ or copies of $v$, the choice of which such edge to add is arbitrary: For every edge $(u,v)$ in $G$ we add the edge $(v_{j},v_{j+1})$ to $E_{C}$ if $v$’s ID is larger than $u$’s, and $(u,v)$ is the $j$’th edge in the clockwise local embedding of $v$.
In [17], $\hat{G}$ was defined without the edges $E_{C}$, as their purposes did not require it. However, the following properties (except for Property 5) of the face-disjoint graph were proven in [17] and are yet valid.

### v4

@@ 6.2 Minimum st-Cut
Given a directed planar graph $G$ with non-negative edge weights and two vertices $s,t$, there is a $\tilde{O}(D^{2})$ round randomized algorithm that finds the minimum $st$-cut w.h.p..
To prove this, we use Theorem 1.2 to compute a maximum $st$-flow $f$ in $G$ in $\tilde{O}(D^{2})$ rounds. By [7], a corresponding minimum $st$-cut is a cut defined by a set of saturated edges (an edge is saturated if the flow it pushes equals its capacity) which are usually found by solving a reachability problem in the residual graph, which for technical reasons we reduce to an SSSP problem that is solved with an exact SSSP algorithm that runs in $\tilde{O}(D^{2})$ rounds for planar graphs by [27]. The cut edges and bisection are then deduced from the distance computation.
Proof.
At first, we use Theorem 1.2 to compute a maximum $st$-flow $f$ in $G$ in $\tilde{O}(D^{2})$ rounds. A minimum $st$-cut is a cut determined by the reachable vertices from $s$ in the residual graph $R$ w.r.t. the flow $f$. This cut is denoted by $\overrightarrow{\delta}(S,V\setminus S)$, where $S$ (respectively $V\setminus S$) is the side of the cut containing $s$ (respectively $t$). The residual graph $R$ is defined as follows. For each edge $(u,v)$ in $G$ with capacity $c(\cdot)$, $R$ has two edges $(u,v)$ and $(v,u)$ with weight c(u,v)f(u,v)c(u,v)-f(u,v) and $f(u,v)$ respectively. Edges of $R$ with weight zero are then removed.
Thus, our goal is to construct $R$ and compute reachability on it. Notice that the removal of zero-weight edges may increase the diameter of the graph. Hence, we instead construct a network $R^{\prime}$ that is similar to $R$ except that in $R^{\prime}$ zero-weight edge (instead of being removed) are assigned a weight of $\infty$ and all other edges are assigned a weight of zero. Then, a vertex $u$ is reachable in $R$ from $s$ iff $\mathsf{dist}(s,u)=0$ in $R^{\prime}$. Note, since the weights of the edges in $R^{\prime}$ are either zero or $\infty$, we can represent $\infty$ by $1$.
We construct $R^{\prime}$ locally (i.e., without communication): since each vertex $v\in G$ knows its neighbors, incident edges, their capacity and amount of flow pushed on them, then it learns locally its incident edges in $^{\prime}. RR^{\prime}$ is embedded in the plane similarly to the embedding of $G$. E.g., the new copy $(v,u)$ of the existing edge $(u,v)$ succeeds it in the local ordering of the higher ID endpoint (say $v$) and precedes it in the local ordering of the other endpoint (say $u$). Note that $R^{\prime}$ might be a multi-graph, however, one round of communication on $R^{\prime}$ can be simulated in two rounds on $G$.
To compute a directed SSSP tree from $s$, we use the exact SSSP algorithm of Li and Parter [27] that runs in $\tilde{O}(D^{2})$ rounds (Theorem 1.6, Lemma 5.3 in their paper). Although it is stated in their paper for the undirected case, it can be easily adapted to the directed case as noted by Parter in [36]. In addition, we note that the same algorithm works for planar multi-graphs as well.
Finally, each $u\in G$ knows whether it is in $S$ or $V\setminus S$ according to its distance from $s$ in $R^{\prime}$. So, every edge $(u,v)$ of $G$ s.t. $\mathsf{dist}(s,u)=0$ and $\mathsf{dist}(s,v)\neq 0$ in $R^{\prime}$ is marked by its endpoints as a cut edge. This is computed in one round.
Next, we show an almost optimal algorithm for $(1+o(1))$ approximation of the undirected minimum $st$-cut, where $s$ and $t$ lie on the same face.
@@
Given an undirected planar graph $G$ with non-negative edge weights and two vertices $s,t$ that lie on the same face of $G$, there is a randomized algorithm that finds an approximate $(1+o(1))$ minimum $st$-cut and runs in $D\cdot n^{o(1)}$ rounds w.h.p..
@@
Concretely, we take a primal-dual detour, where we consider faces in the dual graph that map to the vertices $s$ and $t$ in the primal [^18] The dual of the dual is the primal, i.e, primal vertices map to faces of the dual graph. Moreover, the dual face that is mapped to any vertex $v$ is the (only) face that consists of all darts entering $v$. See [26] for more details.. See Figure 15. We refer to faces of the dual as dual faces. A special case of the cycle-cut duality (3.1) is given by Reif [39], stating: An $st$-separating cycle in the dual graph is the dual of an $st$-cut in the primal, where a cycle in the dual graph is said to be $st$-separating, iff it contains the the dual face that is mapped to $s$ in its interior, and the dual face mapped to $t$ in its exterior (or vice versa). Thus, we find a $(1+\epsilon)$-approximation of the minimum $st$-separating cycle in $G^{*}$ in order to find its dual, the desired $(1+\epsilon)$-approximate minimum $st$-cut in $G$ (then, setting $\epsilon=1/n^{o(1)}$ gives our promised $(1+o(1))$-approximation).
@@
Note, any $st$-cut in $G\cup\{e\}$ shall contain $e$, and removing it gives $G$. Thus, to have the min $st$-cut in $G$, we can look for the min $st$-cut in $G\cup\{e\}$. In particular, there is a mapping between $st$-cuts of $G$ and $G\cup\{e\}$, where the value of the cut in $G\cup\{e\}$ is the same value of the corresponding cut in $G$ plus the weight ($nW$) of $e$. Thus, when applying the duality of Reif we consider $G^{*}_{virt}$. That is, we look for an $st$-separating cycle in $G^{*}_{virt}$. By duality, this cycle should contain $e^{*}=(f_{1},f_{2})$.
@@
Note, the edge $e^{*}=(f_{1},f_{2})$ closes a cycle with $P$, denoted $C=P\cup\{e^{*}\}$, that would define the desired $st$-separating cycle. $C$ is separating as any cycle in a planar graph is separating (Jordan curve theorem [45]), in particular it is $st$-separating, for: (1) $s$ and $t$ are faces in the dual, each has to be either in the interior or the exterior of $C$, (2) one dart of $e$ is in $C$’s interior and the other in its exterior, each such dart corresponds to a distinct dual face, one of which maps to $s$ and the other maps to $t$ because $e$ connects $s$ and $t$ in the primal.
Since $C=P\cup\{e\}$, then the length of $C$ is that of $P$ plus $nW$. Contracting $e$ subtracts $nW$ from the weight of $C$ and results with a $(1+\epsilon)$-approximate $st$-separating cycle in $G^{*}$, for otherwise, $P$ is not a $(1+\epsilon)$ approximation to the shortest $f_{1}$-to-$f_{2}$ path, a contradiction. I.e., any $f_{1}$-to-$f_{2}$ path $P$ in $(G\cup\{e\})^{*}$ maps to a $f$-to-$f$ path $P$ (cycle) in $G^{*}$ of the same weight.
@@
Next, we aim to mark those $st$-cut edges in $G$. By the above, it is sufficient to mark edges in $G$ who are the duals of edges $P$ in $G^{*}_{virt}$. Assuming we already applied Theorem 1.3, in particular, we have computed an approximate SSSP tree $T$ from $f_{1}$, then, we want to mark $P$. We already know that we can simulate any minor-aggregation algorithm on $G^{*}_{virt}$ (for details, one is referred to the proof of Theorem 1.3), thus, we can simulate any minor-aggregation algorithm on any subgraph of $G^{*}_{virt}$ (Corollary 11 of [18]). I.e., we can invoke any minor-aggregation algorithm on $T$. To mark $P$ edges, a minor-aggregation tree-rooting and subtree-sum procedures of [18] are invoked, where we root $T$ with $f_{1}$ and compute subtree-sum on the rooted tree, such that, the input value of all nodes is zero, except for $f_{2}$ that would have an input of one. The edges whose both endpoints have subtree-sum value of one are marked, those exactly are the $P$-edges. Those procedures are of $\tilde{O}(1)$ minor-aggregation rounds, which are simulated within $\tilde{O}(D)$ rounds in $G$ by applying Theorem 4.14, the exact details of this simulation on $G^{*}_{virt}$ are provided in the proof of Theorem 1.3.
@@
Next, we show how to find the cut vertices, in particular, we show how to find a set $S$ such that $P=\delta(S,V\setminus S)$. To do that, we take an approach similar to that described in the proof of Theorem 6.1, where we define a graph $R^{\prime}$, in which, each $P$-edge is assigned a weight of one, and all other edges a weight of zero, we then again use the minor-aggregation SSSP approximation algorithm of [43] with an arbitrary constant approximation parameter $\epsilon$ in order to compute a $(1+\epsilon)$-approximate SSSP tree $T$ from $s$ in $R^{\prime}$. In particular, we modify the algorithm to work with edges with weights in $0,n^{O(1)}]$ rather in $1,n^{O(1)}]$, as shown in the proof of Theorem 1.3. Finally, vertices whose distance is zero from $s$ are in $S$ (other vertices are in $V\setminus S$).
@@
The round complexity of simulating the SSSP approximation algorithm is $D\cdot n^{o(1)}$, since the part-wise aggregation problem can be solved on $G$ in $\tilde{O}(D)$ $\mathsf{CONGEST}$ rounds (Corollary 4.6), and the approximate SSSP algorithm of [43] with $\epsilon=1/n^{o(1)}$ runs in $n^{o(1)}$ minor-aggregation rounds. Thus, by Lemma 4.8, simulating the algorithm incurs a total of $D\cdot n^{o(1)}$ $\mathsf{CONGEST}$ rounds in $G$. Note, after applying the distance approximation algorithm, vertices of $G$ know locally to what side of the cut do they belong. ∎
@@ 7 Directed Global Minimum Cut
Preprocessing: We work with the network obtained from $G$ by adding the reversal dart of each edge to the graph as described in the proof of Theorem 2.1. We abuse notation and denote this network by $G$. The reversal dart gets assigned a weight of zero. It is not hard to see that the new network can be constructed and embedded in the plane without any need for communication (assuming an embedding of the original network). Note that any $\mathsf{CONGEST}$ round on the augmented network is simulated within two rounds on the communication graph.
@@
    $ID(C_{X_{i}})$ for each child bag $X_{i}$ of $X$ is aggregated on $X$ in $\tilde{O}(D)$ rounds so that the ID of the minimum weight cycle $C_{X}$ (that does not cross $F_{\!{}_{X}}$) is known to all vertices in $X$. Note, vertices in $X_{i}$s know $ID(C_{X_{i}})$, thus they can broadcast it (this is obviously true for leaf bags, for non-leaf bags, this is recursively maintained in the next two steps). Since $X$ is of $\tilde{O}(D)$ diameter, this aggregation terminates in $\tilde{O}(D)$ rounds. Next, we consider cycles that cross $F_{\!{}_{X}}$.
@@
      Due to the above claim, we get that $P$ combined with $d$ gives the minimum weight simple dart cycle that uses $d$, denoted $C_{d}$. We define $C$ to be the minimum over $C_{d}$ for all $S_{\!{}_{X}}$ darts $d$. If $C$ is of weight strictly less than $C_{X}$, then $C_{X}$ is set to $C$ and is identified with the ID of an endpoint of $d$ ($d$ is the dart s.t. $C=C_{d})$) concatenated with the weight of $C$.
@@
      $C$ does not use $S_{\!{}_{X}}$: In this case, we iterate over node-parts $g_{i}$ in $F_{\!{}_{X}}$, in each iteration we remove $S_{\!{}_{X}}$ darts from $DDG(f)$ and all clique edges between all node-parts $g_{1},g_{2},\ldots$ that correspond to the same face $g$ of $G$ as $g_{i}$.
      Figure 18: $DDG(f)$ in the iteration of node-part $g_{2}$. $g_{1},g_{2},g_{3}$ correspond to the face $g$ of $G$, where each shadow color demonstrates a distinct child bag of $X^{*}$. Black vertices are endpoints of $S_{\!{}_{X}}$ edges and red vertices are node parts in $X^{*}$’s child bags. All $S_{\!{}_{X}}$ darts are removed from the DDG, as well as clique edges on $g_{1},g_{2},g_{3}$. All other edges have both their darts present (for simplicity, some are undirected in the picture). The green directed path demonstrates a shortest $g_{2}$-to-$g_{1}$ path, representing a cycle that contains $g$ in $G^{*}$.
@@
    Note that $T$ is represented by $\hat{G}$, our tool for aggregations on $G^{*}$. Thus, we next work with it and with minor-aggregations on $G^{*}$ and $T$, which by Section 4 is possible since (1) we can do that on $G^{*}$ or any minor of it, and (2) $T$ is a subgraph (minor) of $G^{*}$. For an extended discussion of $\hat{G}$ and minor-aggregations on $G^{*}$, see Sections 3, 4, and Appendix A.
@@
Appendix A Deferred Proofs from Section 3 (Properties of $\hat{G}$)
We prove the following properties of the face-disjoint graph. All of those properties (except for Property 5) were proven in [17]. However, for completeness a proof is provided for all properties.